Find an equation of the line satisfying the given conditions
Through (-4,3); perpendicular to -8x+5y=77
possible answers:
5x+8y=4
5x-8y=77
-8x-5y=4
5x-8y=4

Answer by Alan3354(69443)

Find an equation of the line satisfying the given conditions
Through (-4,3); perpendicular to -8x+5y=77
------------------
Find the slope of -8x+5y=77 by putting it in slope-intercept form y=mx+b (that means solve for y).
-8x+5y=77
5y = 8x + 77
y = (8/5)x + 77/5
The slope, m, is 8/5. Lines perpendicular will have a slope that's the negative invervse, m = -5/8.
------------------
y-y1 = m*(x-x1) where (x1,y1) is the point (-4,3)
y-3 = (-5/8)*(x+4)
8y-24 = -5x - 20
5x+8y = 4